OpenStudy (anonymous):

Evaluate the integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ x^2\sqrt{12x+1} }dx\]

OpenStudy (anonymous):

I got to the point where I have to use partial fractions.

OpenStudy (anonymous):

\[24\int\limits_{}^{}\frac{ 1 }{(s^2-1)^2 }ds\]

OpenStudy (anonymous):

I am stuck here.

OpenStudy (tkhunny):

Lot's of factoring. (s+1)(s-1) repeated.

OpenStudy (anonymous):

My work: \[\int\limits\limits_{}^{}\frac{ 1 }{ x^2\sqrt{12x+1} }dx\] u=12x+1 x=(u-1)/12 1/12 du=dx \[\int\limits_{}^{}\ \frac{ 1 }{ (\frac{ u-1 }{ 12 })^2\sqrt{u}}du\] \[12\int\limits_{}^{}\frac{ 1 }{ (u-1)^2\sqrt{u} }du\] s=√ (u) s^2=u ds=1/2√ u) du 2√ (u)ds=du \[24\int\limits_{}^{}\frac{ 1 }{ (s^2-1)^2 }ds\]

OpenStudy (anonymous):

I am stuck here.

OpenStudy (anonymous):

I feel that I could integrate that easily but something stupid is preventing me from doing so.

OpenStudy (anonymous):

I am pretty sure I have to use partial fractions but I don't know how...

hartnn (hartnn):

(s^2-1)^2 = (s+1)^2 (s-1)^2 A/(s+1) +B/(s+1)^2 + C/(s-1)+ D/(s-1)^2

OpenStudy (anonymous):

How did you break it up? @hartnn

hartnn (hartnn):

generally, (ax+b)^2 in the denominator is break up like A/(ax+b) +B/(ax+b)^2

OpenStudy (anonymous):

Sorry. I JUST learnt this :P .

hartnn (hartnn):

the numerator should be 1 degree less than the denom for partial fractions. so for 1/(ax+b) it'll be A/(ax+b) and (ax+b)^2 in the denominator is break up like A/(ax+b) +B/(ax+b)^2

OpenStudy (anonymous):

Allright, How did you get this though? A/(s+1) +B/(s+1)^2 + C/(s-1)+ D/(s-1)^2

hartnn (hartnn):

for 1/(s+1)^2 A/(s+1) +B/(s+1)^2 got this ? same thing for 1/(s-1)^2

OpenStudy (anonymous):

A/(s+1) +B/(s+1)^2 Explain that? Sorry, I am simply not understanding the partial fraction decomposition :P .

hartnn (hartnn):

for 1/(x) you would have written A/x right ?

OpenStudy (anonymous):

Yeah, if I didn't know that was 1.

hartnn (hartnn):

for 1/(x+1) ?

OpenStudy (anonymous):

A/(x+1) ?

OpenStudy (anonymous):

No wait.

OpenStudy (anonymous):

B/(x+1)

OpenStudy (anonymous):

Would have a different coefficient.

hartnn (hartnn):

ok, whatever now we go for ^2 in denominator. how would u write 1/x^2 ?

hartnn (hartnn):

(Ax+B)/x^2 or A/x +B/x^2

hartnn (hartnn):

similarly for 1/x^3 (Ax^2+Bx+C)/x^3 or A/x +B/x^2+C/x^3

hartnn (hartnn):

see the pattern ?

OpenStudy (anonymous):

I know that by memorizing but why?

OpenStudy (anonymous):

But yep I see that :) .

hartnn (hartnn):

the numerator should be upto 1 degree less than the denom for partial fractions.

OpenStudy (anonymous):

Yep, that makes sense.

hartnn (hartnn):

k, so would u able to find A,B,C,D now ?

OpenStudy (anonymous):

Ahh, now I see how you split it up.

OpenStudy (anonymous):

Umm.... Maybe, let me do it out.

hartnn (hartnn):

okk..

OpenStudy (anonymous):

@hartnn : I expanded it out and everything but surely, there must be an easier way to solve for A,B,C,D ?

hartnn (hartnn):

to find B put s=-1 to find D, put s=+1

OpenStudy (anonymous):

Because I have a MASSIVE expansion.

OpenStudy (anonymous):

Oh right!

hartnn (hartnn):

to find A,C put any other values of s say,s=0 s=2

OpenStudy (anonymous):

But if I put B=-1 I get 1=0 which is not true.

hartnn (hartnn):

put B =-1 ?

OpenStudy (anonymous):

\[\frac{ 1 }{ (s+1)^2(s-1)^2 }=\frac{ A }{ (s+1) }+\frac{ B }{ (s+1)^2 }+\frac{ C }{ (s-1) }+\frac{ D }{ (s-1)^2 }\] \[1=A(s+1)^2(s-1)^3+B(s+1)(s-1)^3+C(s+1)^3(s-1)^2+D(s+1)^3(s-1)\]

OpenStudy (anonymous):

so if I put s=-1 every term becomes 0.

OpenStudy (anonymous):

or s=1 for that matter.

hartnn (hartnn):

that expansion is incorrect.

OpenStudy (anonymous):

:( .

hartnn (hartnn):

\(1=A(s+1)(s-1)^2+B(s-1)^2+C(s+1)^2(s-1)+D(s+1)^2\)

OpenStudy (anonymous):

Let me try again...

OpenStudy (anonymous):

So I would go A(s+1)^2(s−1)^3 right?

OpenStudy (anonymous):

Because I am multiplying by every other term in the denominator on the RHS.

hartnn (hartnn):

no.... you rae not doing that..

OpenStudy (anonymous):

\[\frac{1}{x^2\sqrt{12x+1}}=\frac{\sqrt{12 x+1}}{x^2}-\frac{12 \sqrt{12 x+1}}{x}+\frac{144}{\sqrt{12 x+1}} \]

OpenStudy (anonymous):

Sigh... I am being stupid.

hartnn (hartnn):

 so you are not multiplying entire bc (here,(s-1)^2(s+1)^2) just the uncommon terms.